Compute the gravitational potential and gravitational field due to a hollow sphere: (i) at a point outside the sphere (ii) at a point inside the hollow space, and (iii) at a point inside the material. 


1
Expert's answer
2021-03-02T18:06:43-0500

i)


"E=\frac{GM}{r^2}, r>R"

ii)


"E=0, r<R"

iii)


"E=\frac{GM}{R^2}, r=R"